# Variational Flow Matching for Graph Generation

### v1

@@ Abstract
We present a formulation of flow matching as variational inference, which we refer to as variational flow matching (VFM). Based on this formulation we develop CatFlow, a flow matching method for categorical data. CatFlow is easy to implement, computationally efficient, and achieves strong results on graph generation tasks. In VFM, the objective is to approximate the posterior probability path, which is a distribution over possible end points of a trajectory. We show that VFM admits both the CatFlow objective and the original flow matching objective as special cases. We also relate VFM to score-based models, in which the dynamics are stochastic rather than deterministic, and derive a bound on the model likelihood based on a reweighted VFM objective. We evaluate CatFlow on one abstract graph generation task and two molecular generation tasks. In all cases, CatFlow exceeds or matches performance of the current state-of-the-art models.
@@ 1 Introduction
In recent years, the field of generative modeling has seen notable advancements. In image generation [35, 38], the development and refinement of diffusion-based approaches — specifically those using denoising score matching [51] — have proven effective for generation at scale [14, 46]. However, while training can be done effectively, the constrained space of sampling probability paths in a diffusion requires tailored techniques to work [44, 56]. This is in contrast to more flexible approaches such as continuous normalizing flows (CNFs) [8], that are able to learn a more general set of probability paths than diffusion models [45], at the expense of being expensive to train as they require one to solve an ODE during each training step (see e.g. [5, 39, 13]).
Recently, Lipman and collaborators [24] proposed flow matching (FM), an efficient and simulation-free approach to training CNFs. Concretely, they use a per-sample interpolation to derive a simpler objective for learning the marginal vector field that generates a desired probability path in a CNF. This formulation provides equivalent gradients without explicit knowledge of the (generally intractable) marginal vector field. This work has been extended to different geometries [7, 20] and various applications [52, 9, 12, 21]. Similar work has been proposed concurrently in [27, 1].
@@
While VFM provides a general formulation, our primary interest in this paper is its application to graph generation, where the data are categorical. This setting leads to a simple method that we refer to as CatFlow, in which the objective reduces to training a classifier over end points on a per-component basis. We apply CatFlow to a set of graph generation tasks, both for abstract graphs [30] and molecular generation [34, 18]. By all metrics, our results match or substantially exceed those obtained by existing methods.
@@ 2.1 Transport Framework for Generative Modeling and CNFs
Common generative modeling approaches such as normalizing flows [36, 32] and diffusion models [14, 46] parameterize a transformation $\varphi$ from some initial tractable probability density $p_{0}$ – typically a standard Gaussian distribution – to the target data density $p_{1}$. In general, there is a trade-off between allowing $\varphi$ to be expressive enough to model the complex transformation while ensuring that the determinant term is still tractable. One such transformation is a continuous normalizing flow (CNF).
@@
In CNFs, this vector field is learned using a neural network $v_{t}^{\theta}$. Through the change of variables formula $p_{t}(x)$ can be evaluated (see appendix D) and hence one could try and optimize the empirical divergence between the resulting distribution $p_{1}$ and target distribution. However, obtaining a gradient sample for the loss requires ones to solve the ODE induced during training, making this approach computationally expensive.
@@ 2.2 Flow Matching
In flow matching [24], the aim is to regress the underlying vector field of a CNF directly on the interval $t\in[0,1]$. Flow matching leverages the fact that even though we do not have access to the actual underlying vector field – which we denote as $u_{t}$ – and probability path $p_{t}$, one can construct a per-example formulation by defining conditional flows, i.e. the trajectories towards specific datapoints $x_{1}$. Concretely, FM sets:
@@ 3 Variational Flow Matching for Graph Generation
We derive CatFlow through a novel, variational view on flow matching we call Variational Flow Matching (VFM). The VFM framework relies on two insights. First, we can define the marginal vector field and its approximation in terms of an expectation with respect to a distribution over end points of the transformation. This implies that we can map a flow matching problem onto a variational counterpart, opening up the usage of variational inference techniques. Second, under typical assumptions on the forward process, we can decompose the expected conditional vector field into components for individual variables which can be computed in terms of the marginals of the distribution over end points of the conditional trajectories. This implies that, without loss of generality, we can solve a VFM problem using a fully-factorized variational approximation, providing a tractable approximate vector field. For categorical data, the corresponding vector field can be computed efficiently via direct summation. This results in a closed-form objective to train CNFs for categorical data, which we refer to as CatFlow. Note that we develop the theory of VFM in section 3 and we relate VFM to flow matching and score-based diffusion in section 4.
@@ 3.1 Flow Matching using Variational Inference
In any flow matching problem, the vector field in eq. 2 can be expressed as an expectation
@@
This observation leads us to propose a change in perspective. Rather than predicting the vector field directly, we can define an approximate vector field in terms of an expectation with respect to a variational distribution $q_{t}^{\theta}$ with parameters $\theta$,
$v_{t}^{\theta}(x):=\int u_{t}(x\mid x_{1})q_{t}^{\theta}(x_{1}\mid x)\mathrm{d}x_{1}.$ (8)
@@
Concretely, we can define a variational flow matching problem by minimizing the Kullback-Leibler (KL) divergence from $p_{t}$ to $q_{t}^{\theta}$, which we can express as (see section A.1 for derivations)
@@
where $t\sim\text{Uniform}(0,1)$ and $x,x_{1}\sim p_{t}(x,x_{1})$.
This leads us to propose the variational flow matching (VFM) objective
@@
Note that, though different, this objective is similar to a typical negative log-likelihood objective.
At this stage of the exposition, it is not yet clear how useful this variational perspective on flow matching is in practice. First, while we can in principle map any flow matching problem onto a variational inference problem, this requires learning an approximation of a potentially complex, high-dimensional distribution $p_{t}(x_{1}\mid x)$. Second, we express $v^{\theta}_{t}(x)$ as an expectation, which could well be intractable. This could mean that we would have to approximate this expectation, for example using a Monte Carlo estimator. We will address these concerns in section 3.2.
@@ 3.2 Mean-Field Variational Flow Matching
At first glance we do not seem to obtain much from this variation view due to the intractability of $p_{t}(x_{1}\mid x)$ and $v_{t}^{\theta}(x)$. Fortunately, we can simplify the objective and the calculation of the marginal vector field under certain conditions on the conditional vector field. Specifically, we will focus on the special case where the conditional vector field $u_{t}(x\mid x_{1})$ is linear in $x_{1}$, such as in straight line interpolations commonly used in flow matching. Informally, we show that in that case the specific form of the distribution used does not affect the expected value of the conditional vector field as long as the marginals coincide. As such, the solution can be found by a mean-field/factorized distribution exactly. This reduces the problem from one high-dimensional problem, into a series of low dimension problems. Formally, the following holds:
@@
We provide a proof in section A.3. It follows directly from theorem 1 that without loss of generality we can consider the considerably easier task of a fully-factorized approximation
@@
Now, to calculate the vector field $v_{t}^{\theta}(x)$, we can simply substitute the factorized distribution $q^{\theta}_{t}(x_{1}\mid x)$ into eq. 8. However, this still requires an evaluation of an expectation. Fortunately, leveraging the linearity condition significantly simplifies this computation. Informally, under this linearity condition, as long as we have access to the first moment of one-dimensional distributions $q^{\theta}_{t}(x_{1}^{d}\mid x)$, we can efficiently calculate $v_{t}^{\theta}(x)$. Therefore, if two distributions $q_{t}(x_{1}\mid x)$ and $q_{t}^{\prime}(x_{1}\mid x)$ share the same first moments, they will describe the same vector field, regardless of any other differences they may have. Note that the training procedure will differ for two distinct distributions – e.g. Gaussian versus Categorical – so the form of the distribution $q^{\theta}_{t}(x_{1}\mid x)$ remains practically important, a flexibility provided through the variational view on flow matching.
Formally, we can rewrite an arbitrary linear conditional vector field as
$u_{t}(x\mid x_{1})=A_{t}(x)x_{1}+b_{t}(x),$ (14)
where $A_{t}(x):[0,1]\times\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}\times\mathbb{R}^{D}$ and $b_{t}(x):[0,1]\times\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$. If we substitute this into the definition of $v_{t}^{\theta}(x)$ in eq. 8, we can use the linearity of the expected value to see that
$v_{t}^{\theta}(x)=\mathbb{E}_{q^{\theta}_{t}(x_{1}\mid x)}\left[A_{t}(x)x_{1}+b_{t}(x)\right]=A_{t}(x)\mathbb{E}_{q^{\theta}_{t}(x_{1}\mid x)}\left[x_{1}\right]+b_{t}(x).$ (15)
@@
Note that this covers both the case of categorical data, which we focus on in this paper, and the case of continuous data, as considered in traditional flow matching methods.
At first glance, the linearity condition of the conditional vector field $u_{t}(x\mid x_{1})$ in theorem 1 might seem restricting. However, in most state-of-the-art generative modeling techniques, this condition is satisfied, e.g. diffusion-based models, such as flow matching [24, 2, 27], diffusion models [14, 46], and models that combine the injection of Gaussian noise with blurring [37, 15], among others [10, 43].
Figure 1: Reparameterized velocities by CatFlow. Note that as $\theta_{t}(x_{t})$ lies in the probability simplex, indeed all velocity implicitly are given the inductive bias of point towards it.
@@ 3.3 CatFlow: Mean-Field Variational Flow Matching for Categorical Data
In CatFlow, we directly apply the VFM framework to the categorical case. Let our parameterised variational distribution $q^{\theta}_{t}(x_{1}^{d}\mid x)=\mathsf{Cat}(x^{d}_{1}\mid\theta_{t}^{d}(x))$, and let us denote $\theta_{t}^{dk}(x):=q_{t}(x_{1}^{d}=k\mid x)$ for brevity. Then, the $d$th component of the learned vector field is
$v_{t}^{\theta,d}(x):=\sum_{k=1}^{K_{d}}\theta^{dk}_{t}(x)\frac{\mathbb{I}[x_{1}^{d}=k]-x}{1-t}$ (17)
@@
$\log q_{t}^{\theta}(x_{1}^{d}\mid x)=\log\prod_{k=1}^{K^{d}}(\theta_{t}^{dk}(x))^{\mathbb{I}[x_{1}^{d}=k]}=\sum_{d=1}^{D}\mathbb{I}[x_{1}^{d}=k]\log\theta_{t}^{dk}(x).$ (18)
@@
$\mathcal{L}_{\text{CatFlow}}(\theta)=-\mathbb{E}_{t,x,x_{1}}\left[\sum_{d=1}^{D}\sum_{k=1}^{K^{d}}\mathbb{I}[x_{1}^{d}=k]\log\theta^{dk}_{t}(x)\right].$ (19)
@@
$\mathbb{E}_{q_{t}^{\theta}(x_{1}^{d}\mid x)}\left[u_{t}(x^{d}\mid x_{1}^{d})\right]=\mathbb{E}_{q_{t}^{\theta}(x_{1}^{d}\mid x)}\left[\frac{x_{1}^{d}-x^{d}}{1-t}\right]=\frac{\theta^{d}(x)-x^{d}}{1-t},$ (20)
since $\theta^{d}(x):=\mathbb{E}_{q_{t}^{\theta}(x_{1}^{d}\mid x)}[u_{t}^{d}(x_{1}^{d}\mid x)]$ and the other terms are not in the expectation. Note that this geometrically corresponds to learning the mapping to a point in the probability simplex, and then flowing towards that. This procedure is illustrated in fig. 1. Because of this, training CatFlow is no less efficient than flow matching.
More precisely, training CatFlow offers three key benefits over standard flow matching. First, the added inductive bias ensures generative paths align with realistic trajectories, improving performance and convergence by avoiding misaligned paths. Second, using cross-entropy loss instead of mean-squared error improves gradient behavior during training, enhancing learning dynamics and speeding up convergence. Lastly, CatFlow’s ability to learn probability vectors, rather than directly choosing classes, allows it to express uncertainty about variables at a specific time. This is especially useful in complex domains like molecular generation, where initial uncertainty about components decreases as more structure is established, leading to more precise predictions.
@@
Graphs, defined by vertices and edges, lack a natural vertex order unlike other data types. This permutation invariance means any vertex labeling represents the same graph if the connections remain unchanged. Note that even though the following results apply to graphs, an unordered set of categorical variables can be described by a graph without edges. Under natural conditions – see appendix A – we ensure CatFlow adheres to this symmetry (see section A.4 for the proof).

### v2

@@ Abstract
We present a formulation of flow matching as variational inference, which we refer to as variational flow matching (VFM). Based on this formulation we develop CatFlow, a flow matching method for categorical data. CatFlow is easy to implement, computationally efficient, and achieves strong results on graph generation tasks. The key observation in VFM is that we can parameterize the vector field of the flow in terms of a variational approximation of the posterior probability path, which is the distribution over possible end points of a trajectory. We show that this variational interpretation admits both the CatFlow objective and the original flow matching objective as special cases. We also relate VFM to score-based models, in which the dynamics are stochastic rather than deterministic, and derive a bound on the model likelihood based on a reweighted VFM objective. We evaluate CatFlow on one abstract graph generation task and two molecular generation tasks. In all cases, CatFlow exceeds or matches performance of the current state-of-the-art.
@@ 1 Introduction
In recent years, the field of generative modeling has seen notable advancements. In image generation ramesh2022hierarchical ; rombach2022high , the development and refinement of diffusion-based approaches — specifically those using denoising score matching vincent2011connection — have proven effective for generation at scale ho2020denoising ; song2020score . However, while training can be done effectively, the constrained space of sampling probability paths in a diffusion requires tailored techniques to work song2020denoising ; zhang2022fast . This is in contrast to more flexible approaches such as continuous normalizing flows (CNFs) chen2018neural , that are able to learn a more general set of probability paths than diffusion models song2021maximum , at the expense of being expensive to train as they require one to solve an ODE during each training step (see e.g. ben2022matching ; rozen2021moser ; grathwohl2018scalable ).
Recently, Lipman and collaborators lipman2023flow proposed flow matching (FM), an efficient and simulation-free approach to training CNFs. Concretely, they use a per-sample interpolation to derive a simpler objective for learning the marginal vector field that generates a desired probability path in a CNF. This formulation provides equivalent gradients without explicit knowledge of the (generally intractable) marginal vector field. This work has been extended to different geometries chen2024flow ; klein2023equivariant and various applications wildberger2024flow ; dao2023flow ; ebrahimy2010stock ; kohler2023flow . Similar work has been proposed concurrently in liu2022flow ; albergo2023stochastic .
@@
While VFM provides a general formulation, our primary interest in this paper is its application to graph generation, where the data are categorical. This setting leads to a simple method that we refer to as CatFlow, in which the objective reduces to training a classifier over end points on a per-component basis. We apply CatFlow to a set of graph generation tasks, both for abstract graphs martinkus2022spectre and molecular generation ramakrishnan2014quantum ; irwin2012zinc . By all metrics, our results match or substantially exceed those obtained by existing methods.
@@ 2.1 Transport Framework for Generative Modeling and CNFs
Common generative modeling approaches such as normalizing flows rezende2015variational ; papamakarios2021normalizing and diffusion models ho2020denoising ; song2020score parameterize a transformation $\varphi$ from some initial tractable probability density $p_{0}$ – typically a standard Gaussian distribution – to the target data density $p_{1}$. In general, there is a trade-off between allowing $\varphi$ to be expressive enough to model the complex transformation while ensuring that the determinant term is still tractable. One such transformation is a continuous normalizing flow (CNF).
@@
In CNFs, this vector field is learned using a neural network $v_{t}^{\theta}$. Through the change of variables formula $p_{t}(x)$ can be evaluated (see appendix˜E) and hence one could try and optimize the empirical divergence between the resulting distribution $p_{1}$ and target distribution. However, obtaining a gradient sample for the loss requires ones to solve the ODE induced during training, making this approach computationally expensive.
@@ 2.2 Flow Matching
In flow matching lipman2023flow , the aim is to regress the underlying vector field of a CNF directly on the interval $t\in[0,1]$. Flow matching leverages the fact that even though we do not have access to the actual underlying vector field – which we denote as $u_{t}$ – and probability path $p_{t}$, one can construct a per-example formulation by defining conditional flows, i.e. the trajectories towards specific datapoints $x_{1}$. Concretely, FM sets:
@@ 3 Variational Flow Matching for Graph Generation
We derive CatFlow through a novel, variational view on flow matching we call Variational Flow Matching (VFM). The VFM framework relies on two insights. First, we can define the marginal vector field and its approximation in terms of an expectation with respect to a distribution over end points of the transformation. This implies that we can map a flow matching problem onto a variational counterpart. Second, under typical assumptions on the forward process, we can decompose the expected conditional vector field into components for individual variables which can be computed in terms of the marginals of the distribution over end points of the conditional trajectories. This implies that, without loss of generality, we can solve a VFM problem using a fully-factorized variational approximation, providing a tractable approximate vector field. For categorical data, the corresponding vector field can be computed efficiently via direct summation. This results in a closed-form objective to train CNFs for categorical data, which we refer to as CatFlow. We develop the theory of VFM in section˜3 and we relate VFM to flow matching and score-based diffusion in section˜4.
@@ 3.1 Flow Matching using Variational Inference
In any flow matching problem, the vector field in eq.˜2 can be expressed as an expectation
@@
This observation leads us to propose a change in parameterization of the learned vector field. Rather than predicting the components of the vector field directly, we can define an approximate vector field in terms of an expectation with respect to a variational distribution $q_{t}^{\theta}$ with parameters $\theta$,
$v_{t}^{\theta}(x):=\int u_{t}(x\mid x_{1})\>q_{t}^{\theta}(x_{1}\mid x)\>\mathrm{d}x_{1}.$ (8)
@@
Concretely, we can define a variational flow matching problem by minimizing the Kullback-Leibler (KL) divergence from $p_{t}$ to $q_{t}^{\theta}$, which we can express as
@@
where $t\sim\text{Uniform}(0,1)$ and $x,x_{1}\sim p_{t}(x,x_{1})$ (see section˜A.1 for derivations). This leads us to propose the variational flow matching (VFM) objective
@@
While this variational formulation of flow matching is promising, two potential drawbacks emerge in practical applications. First, although it is feasible to reformulate any flow matching problem as a variational inference task, doing so requires learning an approximation of a potentially complex, high-dimensional distribution $p_{t}(x_{1}\mid x)$ (including any correlations between the different $x_{1}^{d}$). Second, representing $v^{\theta}_{t}(x)$ as an expectation may pose intractability challenges. Interestingly, under typical assumptions about the conditional velocity field in flow matching, this objective simplifies significantly, making it no more computationally demanding than standard flow matching, a point we will further address in section˜3.2.
@@ 3.2 Mean-Field Variational Flow Matching
At first glance, we do not seem to obtain much from this variational view due to the intractability of $p_{t}(x_{1}\mid x)$ and $v_{t}^{\theta}(x)$. Fortunately, we can simplify the objective and the calculation of the marginal vector field under the typical case where the conditional vector field $u_{t}(x\mid x_{1})$ is linear in $x_{1}$, such as in straight line interpolations commonly used in flow matching. This leads to two simplifications. First, we notice that the expected value of a linear conditional velocity field simply equals the conditional velocity field towards the expectation/mean of the distribution we are considering, i.e.
$u_{t}(x)=\mathbb{E}_{p_{t}(x_{1}\mid x)}\left[u_{t}(x\mid x_{1})\right]=u_{t}(x\mid\mathbb{E}_{p_{t}}\left[x_{1}\mid x\right])\text{ if $u_{t}(x\mid x_{1})$ is linear in $x_{1}$}.$ (11)
This means that as long as our variational distribution has the same mean as $p_{t}(x_{1}\mid x)$, we will learn the same underlying field. Second, we realize that the expected value of $x_{1}^{d}$ only depends on $p_{t}(x_{1}^{d}\mid x)$, and thus as such as long as $q_{t}^{\theta}(x_{1}\mid x)$ has the same marginal expectations $\mathbb{E}\left[x_{1}^{d}\mid x\right]$, we will learn the same field as under $p_{t}(x_{1}\mid x)$. In other words, it suffices to learn a fully-factorized approximation $q_{t}^{\theta}(x_{1}\mid x)$, there is no need to fully characterize the covariance of the posterior probability path. Thich allows us to reduce a high-dimensional variational problem to a series of low dimension problems.
Formally, the following holds:
@@
We provide a proof in section˜A.3.
It follows directly from theorem˜1 that without loss of generality we can consider the considerably easier task of a fully-factorized approximation
@@
Even though we use a factorized model, it is worth emphasizing that due to the linearity of the conditional field, a mean-field variational distribution can learn the solution exactly.
@@
To calculate the vector field $v_{t}^{\theta}(x)$, we can simply substitute the factorized distribution $q^{\theta}_{t}(x_{1}\mid x)$ into eq.˜8. However, this still requires an evaluation of an expectation. Fortunately, leveraging the linearity condition significantly simplifies this computation, since as long as we have access to the first moment of one-dimensional distributions $q^{\theta}_{t}(x_{1}^{d}\mid x)$, we can efficiently calculate $v_{t}^{\theta}(x)$ by simply considering the conditional field towards it. Note that for many distributions – e.g. Gaussian – learning its parameters is equivalent to learning its expected value. Note that the training procedure will differ for two distinct distributions – e.g. Gaussian versus Categorical – so the form of the distribution $q^{\theta}_{t}(x_{1}\mid x)$ remains practically important, a flexibility provided through the variational view on flow matching.
@@
Note that this covers both the case of categorical data, which we focus on in this paper, and the case of continuous data, as considered in traditional flow matching methods. We provided the general algorithm for training and generation in section˜B.1.
At first glance, the linearity condition of the conditional vector field $u_{t}(x\mid x_{1})$ in theorem˜1 might seem restricting. However, in most state-of-the-art generative modeling techniques, this condition is satisfied, e.g. diffusion-based models, such as flow matching lipman2023flow ; albergo2022building ; liu2022flow , diffusion models ho2020denoising ; song2020score , and models that combine the injection of Gaussian noise with blurring rissanen2022generative ; hoogeboom2022blurring , among others daras2022soft ; singhal2023diffuse .
Figure 1: Parameterization of the vector field in CatFlow. Given an interpolant $x_{t}=tx_{1}+(1-t)x_{0}$, CatFlow predicts a categorical distribution $q^{\theta}_{t}(x_{1}\mid x_{t})$ parameterized by a vector $\mu_{t}(x_{t})$. The resulting construction for the vector field $v_{t}^{\theta}(x_{t})=(\mu_{t}(x_{t})-x_{t})/(1-t)$ ensures that trajectories converge to a point on the simplex at $t=1$.
@@ 3.3 CatFlow: Mean-Field Variational Flow Matching for Categorical Data
In CatFlow, we directly apply the VFM framework to the categorical case. Let our parameterised variational distribution $q^{\theta}_{t}(x_{1}^{d}\mid x)=\mathsf{Cat}(x^{d}_{1}\mid\theta_{t}^{d}(x))$, and let us denote the parameters of this categorical as
$\mu_{t}^{dk}(x):=q^{\theta}_{t}(x_{1}^{d}=k\mid x)=\mathbb{E}_{q^{\theta}_{t}(x_{1}\mid x)}[\mathbb{I}[x_{1}^{d}=k]].$ (16)
Then, the $d$th component of the learned vector field is
$v_{t}^{\theta,d}(x):=\sum_{k=1}^{K_{d}}\mu^{dk}_{t}(x)\frac{\mathbb{I}[x_{1}^{d}=k]-x}{1-t}.$ (17)
@@
$\log q_{t}^{\theta}(x_{1}^{d}\mid x)=\log\prod_{k=1}^{K^{d}}(\mu_{t}^{dk}(x))^{\mathbb{I}[x_{1}^{d}=k]}=\sum_{d=1}^{D}\mathbb{I}[x_{1}^{d}=k]\log\mu_{t}^{dk}(x).$ (18)
@@
$\mathcal{L}_{\text{CatFlow}}(\theta)=-\mathbb{E}_{t,x,x_{1}}\left[\sum_{d=1}^{D}\sum_{k=1}^{K^{d}}\mathbb{I}[x_{1}^{d}=k]\log\mu^{dk}_{t}(x)\right].$ (19)
@@
$\mathbb{E}_{q_{t}^{\theta}(x_{1}^{d}\mid x)}\left[u_{t}(x^{d}\mid x_{1}^{d})\right]=\mathbb{E}_{q_{t}^{\theta}(x_{1}^{d}\mid x)}\left[\frac{x_{1}^{d}-x^{d}}{1-t}\right]=\frac{\mu^{d}(x)-x^{d}}{1-t},$ (20)
since $\mu^{d}(x):=\mathbb{E}_{q_{t}^{\theta}(x_{1}^{d}\mid x)}[u_{t}^{d}(x_{1}^{d}\mid x)]$ and the other terms are not in the expectation. Note that this geometrically corresponds to learning the mapping to a point in the probability simplex, and then flowing towards that. This procedure is illustrated in fig.˜1. Because of this, training CatFlow is no less efficient than flow matching. We provided the algorithm for training and generation for CatFlow and a Gaussian VFM objective in section˜B.2 and section˜B.3 respectively.
More precisely, training CatFlow offers two key benefits over standard flow matching. First, given that CatFlow predicts points in the probability simplex and parametrizes the velocities to point into it, an inductive bias is introduced ensuring all generative paths align with realistic trajectories, hence avoiding misaligned paths. Second, using a cross-entropy loss instead of a mean-squared error improves gradient behavior during training. Both aspects enhance learning dynamics and speed up convergence, which is evaluated in section˜6.3. Lastly, CatFlow’s ability to learn probability vectors – rather than directly choosing classes as is common in discrete approaches – allows the model to express uncertainty about variables at a specific time. This is especially useful in complex domains like molecular generation, where initial uncertainty about components decreases as more structure is established, leading to more precise predictions.
@@
Graphs, defined by vertices and edges, lack a natural vertex order unlike other data types. This permutation invariance means any vertex labeling represents the same graph if the connections remain unchanged. Note that even though the following results apply to graphs, an unordered set of categorical variables can be described by a graph without edges. Under natural conditions – see appendix˜A – we ensure CatFlow adheres to this symmetry (see section˜A.4 for the proof).